# Singular-degenerate parabolic systems with the conormal boundary condition on the upper half space

### v1

@@ 2.2. Main theorems
We impose the following partially BMO regularity assumption on the coefficents $A^{ij}$ and $C_{0}$.
@@
For any $r\in(0,R_{0}]$ and $z_{0}=(z_{0}^{\prime},x_{0d})\in\overline{\Omega_{T}}$, we have
$\max_{1\leq i,j\leq d}\fint_{Q_{r}^{+}(z_{0})}|A^{ij}(t,x)-[A^{ij}]_{r,z_{0}^{\prime}}(x_{d})|d\mu(z)+\fint_{Q_{r}^{+}(z_{0})}|C_{0}(t,x)-[C_{0}]_{r,z_{0}^{\prime}}(x_{d})|d\mu(z)\leq\gamma_{0},$
where $d\mu(z)=x_{d}^{\alpha}dxdt$, $A^{ij}]_{r,z_{0}^{\prime}}(x_{d})$ is the average of $A^{ij}$ with respect to $(t,x^{\prime})$ in $Q_{r}^{\prime}(z_{0}^{\prime})$:
$A^{ij}]_{r,z_{0}^{\prime}}(x_{d})=\fint_{Q_{r}^{\prime}(z_{0}^{\prime})}A^{ij}(t,x^{\prime},x_{d})dx^{\prime}dt,\quad[C_{0}]_{r,z_{0}^{\prime}}(x_{d})=\fint_{Q_{r}^{\prime}(z_{0}^{\prime})}C_{0}(t,x^{\prime},x_{d})dx^{\prime}dt.$
@@
Let $\alpha\in(-1,\infty),p,q\in(1,\infty),\kappa\in(0,1],R_{0}\in(0,\infty)$, and $w=w_{0}(t)w_{1}(x)$, where $w_{0}\in A_{q}(\mathbb{R})$ and $w_{1}\in A_{p}(\mathbb{R}^{d}_{+},\mu)$. Then, there exist $\gamma_{0}=\gamma_{0}(d,n,\alpha,p,q,\kappa,K)\in(0,1)$ and $\lambda_{0}=\lambda_{0}(d,n,\alpha,p,q,\kappa,K,K_{1},R_{0})>0$ such that the following assertions hold. Suppose that (2.1), (2.2), and Assumption 2.1 $(\gamma_{0},R_{0})$ are satisfied.
@@
    Suppose that $|B|\leq K_{1}x_{d}$, $|\hat{B}|\leq K_{1}x_{d}$, and $|C|\leq K_{1}x_{d}^{2}$. For any $F\in L_{q,p}(\Omega_{T},wd\mu)^{d+1}$ with $F_{0}=0$, $f\in L_{q,p}(\Omega_{T},wd\mu)$, and $\lambda\geq\lambda_{0}$, there exists a unique weak solution $u\in\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu)$ to (1.1)-(1.2). Moreover, $u$ satisfies
    $\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+N\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}$ (2.5)
    for some $N=N(d,n,\alpha,p,q,\kappa,K)>0$.
@@
    If additionally, $\alpha\in(0,\infty)$ and $w_{1}(x)=w_{2}(x^{\prime})w_{3}(x_{d})$ with $w_{2}\in A_{p}(\mathbb{R}^{d-1})$ and $w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap X_{p}(\mu)$, then the result of (i) holds with
    $|B|\leq\max(\varepsilon,K_{1}x_{d}),\quad|\hat{B}|\leq K_{1}x_{d},\quad\text{and}\quad|C|\leq\max(K_{1}x_{d},K_{1}x_{d}^{2})$
    if $\varepsilon=\varepsilon(d,n,\alpha,p,q,\kappa,K)\in(0,1)$ is small enough. Moreover,
    $\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}+\|u/x_{d}\|_{L_{q,p}(\Omega_{T},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}\\ \leq N\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+N\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}.$ (2.6)
@@
Let $\alpha\in(0,\infty),p,q\in(1,\infty)$, $\kappa\in(0,1]$, $R_{0}\in(0,\infty)$, $w=w_{0}(t)w_{2}(x^{\prime})w_{3}(x_{d})$, where $w_{0}\in A_{q}(\mathbb{R}),w_{2}\in A_{p}(\mathbb{R}^{d-1})$, and $w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap Y_{p}(\mu)$. Then, there exists $\gamma_{0}=\gamma_{0}(d,n,\alpha,p,q,\kappa,K)\in(0,1),\lambda_{0}=\lambda_{0}(d,n,\alpha,p,q,\kappa,K,K_{1},R_{0})>0$, and $\varepsilon=\varepsilon(d,n,\alpha,p,q,\kappa,K)\in(0,1)$ such that the following assertions hold. Suppose that (2.1), (2.2), and Assumption 2.1 $(\gamma_{0},R_{0})$ are satisfied.
@@
    Suppose that $|B|\leq K_{1}x_{d}$, $|\hat{B}|\leq\max(\varepsilon,K_{1}x_{d})$, and $|C|\leq\max(K_{1}x_{d},K_{1}x_{d}^{2})$. For any $F\in L_{q,p}(\Omega_{T},wd\mu)^{d+1}$, $f\in L_{q,p}(\Omega_{T},wd\mu)$, and $\lambda\geq\lambda_{0}$, there exists a unique weak solution $u\in\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu)$ to (1.1)-(1.2). Moreover, estimate (2.5) holds.
@@
    If additionally, $\alpha\in(1,\infty)$ and $w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap X_{p}(\mu)\cap Y_{p}(\mu)$, then the result of (i) holds with $|B|\leq\max(\varepsilon,K_{1}x_{d})$, $|\hat{B}|\leq\max(\varepsilon,K_{1}x_{d})$, $|C|\leq\max(\varepsilon,K_{1}x_{d},K_{1}x_{d}^{2})$, and (2.6) in place of (2.5).
@@
    The classes $X_{p}(\mu)$ and $Y_{p}(\mu)$ are special cases of general classes of weights introduced by Muckenhoupt in [40], in which the author obtained a complete characterization of weights admitting certain weighted Hardy’s inequalities on a half-line. The class of $X_{p}(\mu)$ weights is characterized by the validity of the weighted Hardy’s inequality $\|u/x_{d}\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N\|D_{d}u\|_{L_{q,p}(\Omega_{T},wd\mu)}$. See Lemma 6.1 for a precise statement. We use this inequality in the proofs of Theorem 2.2 (ii), Theorem 2.3 (ii), and estimate (2.6).
@@
    In Theorem 2.3, from the $Y_{p}(\mu)$ condition it also follows that any weak solution $u$ to (1.1) with $u,Du,F,f\in L_{q,p}(\Omega_{T},wd\mu)$ automatically satisfies the conormal boundary condition (1.2). See the end of Section 6 for the proof. This fact is of independent interest, we will not however use it later in the proofs.
@@
Here $A^{ij},C_{0},B^{i},\hat{B}^{i},C:\mathbb{R}^{d}_{+}\to\mathbb{R}^{n\times n}$ and $u,F^{i},F_{0},f:\mathbb{R}^{d}_{+}\to\mathbb{R}^{n\times n}$ are functions of $x\in\mathbb{R}^{d}_{+}$. We say that $u$ is a weak solution to (2.7)-(2.8) if
@@
holds for all $\varphi\in C_{0}^{\infty}(\mathbb{R}^{d})$. We only state the following elliptic analogue of Theorem 2.2. The elliptic analogue of Theorem 2.3 can be stated similarly.
@@
Let $\alpha\in(-1,\infty),p\in(1,\infty)$, $\kappa\in(0,1]$, $R_{0}\in(0,\infty)$, and $w=w(x)\in A_{p}(\mathbb{R}^{d}_{+})$. Then, there exist $\gamma_{0}=\gamma_{0}(d,n,\alpha,p,q,\kappa,K)\in(0,1)$ and $\lambda_{0}=\lambda_{0}(d,n,\alpha,p,q,\kappa,K,K_{1},R_{0})>0$ such that the following assertions hold. Suppose that (2.1), (2.2), and Assumption 2.1 $(\gamma_{0},R_{0})$ are satisfied.
@@
    Suppose that $|B|\leq K_{1}x_{d}$, $|\hat{B}|\leq K_{1}x_{d}$, and $|C|\leq K_{1}x_{d}^{2}$. For any $F\in L_{p}(\mathbb{R}^{d}_{+},wd\mu)^{d+1}$ with $F_{0}=0$, $f\in L_{p}(\mathbb{R}^{d}_{+},wd\mu)$, and $\lambda\geq\lambda_{0}$, there exists a unique weak solution $u\in W^{1}_{p}(\mathbb{R}^{d}_{+},wd\mu)$ to (2.7)-(2.8). Moreover, $u$ satisfies
    $\|Du\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}\leq N\|F\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+N\|f\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}$ (2.9)
    for some $N=N(d,n,\alpha,p,q,\kappa,K)>0$.
@@
    If additionally $\alpha\in(0,\infty)$ and $w(x)=w_{2}(x^{\prime})w_{3}(x_{d})$ with $w_{2}\in A_{p}(\mathbb{R}^{d-1})$ and $w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap X_{p}(\mu)$, then the result of (i) holds with
    $|B|\leq\max(\varepsilon,K_{1}x_{d}),\quad|\hat{B}|\leq K_{1}x_{d},\quad\text{and}\quad|C|\leq\max(K_{1}x_{d},K_{1}x_{d}^{2})$
    if $\varepsilon=\varepsilon(d,n,\alpha,p,q,\kappa,K)\in(0,1)$ is small enough. Moreover,
    $\|Du\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+\|u/x_{d}\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}\\ \leq N\|F\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+N\|f\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}.$ (2.10)
Theorem 2.5 can be derived from Theorem 2.2 by viewing solutions to elliptic systems as steady state solutions of the corresponding parabolic systems. More precisely, a-priori estimates (2.9) and (2.10) can be derived from estimates (2.5) and (2.6) in Theorem 2.2. See, for example, the proofs of [36, Theorem 2.6] or [16, Theorem 1.2] for details. The existence of solutions can be obtained by following the proof of existence in Theorem 2.2. We therefore omit the proof of Theorem 2.5.

### v2

@@ 2.2. Main theorems
We impose the following partially BMO regularity assumption on the coefficents $A^{ij}$ and $C_{0}$. Let $\gamma_{0}>0$ and $R_{0}\in(0,\infty]$.
@@
For any $r\in(0,R_{0}]$ ($r\in(0,\infty)$ if $R_{0}=\infty$) and $z_{0}=(z_{0}^{\prime},x_{0d})\in\overline{\Omega_{T}}$, there exist measurable $A^{ij}]_{r,z_{0}}(x_{d}),[C_{0}]_{r,z_{0}}(x_{d}):\mathbb{R}_{+}\to\mathbb{R}^{n\times n}$, $1\leq i,j\leq d$, satisfying (2.1)-(2.2) for a.e. $x_{d}\in\mathbb{R}_{+}$, and
$\fint_{Q_{r}^{+}(z_{0})}|A(t,x)-[A]_{r,z_{0}}(x_{d})|d\mu(z)\leq\gamma_{0},\quad\fint_{Q_{r}^{+}(z_{0})}|C_{0}(t,x)-[C_{0}]_{r,z_{0}}(x_{d})|d\mu(z)\leq\gamma_{0},$
where $d\mu(z)=x_{d}^{\alpha}dxdt$.
@@
Let $\alpha\in(-1,\infty),p,q\in(1,\infty),\kappa\in(0,1],K\in[1,\infty),R_{0}\in(0,\infty]$, and $K_{1}\in[0,\infty)$. Let $P=(d,\alpha,p,q,\kappa,K)$ denote the collection of listed parameters. There exist $\gamma_{0}=\gamma_{0}(P)>0$ and $\lambda_{0}=\lambda_{0}(P)>0$ such that the following assertions hold. Let $w=w_{0}(t)w_{1}(x)$, where $w_{0}\in A_{q}(\mathbb{R}),w_{1}\in A_{p}(\mathbb{R}^{d}_{+},\mu)$ with $w_{0}]_{A_{q}(\mathbb{R})}\leq K,[w_{1}]_{A_{p}(\mathbb{R}^{d}_{+},\mu)}\leq K$. Suppose that (2.1), (2.2), and Assumption 2.1 $(\gamma_{0},R_{0})$ are satisfied. Also assume that $F_{0}=0$.
@@
    Suppose that $|B|\leq K_{1}x_{d}$, $|\hat{B}|\leq K_{1}x_{d}$, and $|C|\leq K_{1}^{2}x_{d}^{2}$. There exist $M=M(P)>0$ and $N=N(P)>0$ such that for any $F=(F^{1},\dots,F^{d})\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})^{d}$, $f\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$, $u\in\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$, and $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$ satisfying (1.1)-(1.2), it holds that
    $\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+N\|f\|_{L_{q,p}(\Omega_{T},wd\mu)},$ (2.5)
    Moreover, for any $F\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})^{d}$, $f\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$, and $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$ with $\lambda>0$, there exists a unique solution $u\in\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$ to (1.1)-(1.2).
@@
    Assume, additionally, that $\alpha\in(0,\infty)$, and $w_{1}(x)=w_{2}(x^{\prime})w_{3}(x_{d})$, where $w_{2}\in A_{p}(\mathbb{R}^{d-1}),w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap X_{p}(\mu)$ with $w_{2}]_{A_{p}(\mathbb{R}^{d-1})}\cdot[w_{3}]_{A_{p}(\mathbb{R}_{+},\mu)}\leq K,[w_{3}]_{X_{p}(\mu)}\leq K_{2},K_{2}\in(0,\infty)$. There exist $\varepsilon=\varepsilon(P)>0,M=M(P)>0$, and $N=N(P)>0$, such that the result of (i) holds under the assumption
    $|B|\leq\max(\varepsilon K_{2}^{-1},K_{1}x_{d}),\quad|\hat{B}|\leq K_{1}x_{d},\quad\text{and}\quad|C|\leq\max(K_{1}K_{2}^{-1}x_{d},K_{1}^{2}x_{d}^{2}).$
    Moreover, in place of (2.5) holds the estimate
    $\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}+K_{2}^{-1}\|u/x_{d}\|_{L_{q,p}(\Omega_{T},wd\mu)}$ $+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}$ (2.6)
    $\leq N\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+N\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}.$
@@
Let $\alpha\in(0,\infty),p,q\in(1,\infty),\kappa\in(0,1],K\in[1,\infty),R_{0}\in(0,\infty],K_{1}\in[0,\infty)$. Let $P=(d,\alpha,p,q,\kappa,K)$ denote the collection of listed parameters. Let $\gamma_{0}=\gamma_{0}(P)>0$ and $\lambda_{0}=\lambda_{0}(P)>0$ be from Theorem 2.2. Let $w=w_{0}(t)w_{2}(x^{\prime})w_{3}(x_{d})$, where $w_{0}\in A_{q}(\mathbb{R}),w_{2}\in A_{p}(\mathbb{R}^{d-1}),w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap Y_{p}(\mu)$ with $w_{0}]_{A_{q}(\mathbb{R})}\leq K,[w_{2}]_{A_{p}(\mathbb{R}^{d-1})}\cdot[w_{3}]_{A_{p}(\mathbb{R}_{+},\mu)}\leq K,[w_{3}]_{Y_{p}(\mu)}\leq K_{3},K_{3}\in(0,\infty)$. Suppose that (2.1), (2.2), and Assumption 2.1 $(\gamma_{0},R_{0})$ are satisfied.
@@
    There exist $\varepsilon=\varepsilon(P)>0,M=M(P)>0$, and $N=N(P)>0$ such that the following assertions holds. Suppose that $|B|\leq K_{1}x_{d}$, $|\hat{B}|\leq\max(\varepsilon K_{3}^{-1},K_{1}x_{d})$, and $|C|\leq\max(K_{1}K_{3}^{-1}x_{d},K_{1}^{2}x_{d}^{2})$. Then for any $F=(F^{1},\dots,F^{d})\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})^{d}$, $F_{0},f\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$, $u\in\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$, and $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$ satisfying (1.1)-(1.2), it holds that
    $\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}$ (2.7)
    $\leq N\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+N\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}+NK_{3}\|F_{0}\|_{L_{q,p}(\Omega_{T},wd\mu)}.$
    Moreover, for any $F\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})^{d}$, $F_{0},f\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$, and $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$ with $\lambda>0$, there exists a unique solution $u\in\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$ to (1.1)-(1.2).
@@
    Assume, additionally, that $\alpha\in(1,\infty)$ and $w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap X_{p}(\mu)\cap Y_{p}(\mu)$ with $w_{3}]_{X_{p}(\mu)}\leq K_{2},K_{2}\in(0,\infty)$. There exist $\varepsilon=\varepsilon(P)>0,M=M(P)>0$, and $N=N(P)>0$ , such that the result of (i) holds under the assumption
    $|B|\leq\max(\varepsilon K_{2}^{-1},K_{1}x_{d}),|\hat{B}|\leq\max(\varepsilon K_{3}^{-1},K_{1}x_{d}),|C|\leq\max(\varepsilon(K_{2}K_{3})^{-1},K_{1}K_{2}^{-1}x_{d},K_{1}K_{3}^{-1}x_{d},K_{1}^{2}x_{d}^{2}).$
    Moreover, in place of (2.7) holds the estimate
    $\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}$ $+K_{2}^{-1}\|u/x_{d}\|_{L_{q,p}(\Omega_{T},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}$ (2.8)
    $\leq N\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+N\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}+NK_{3}\|F_{0}\|_{L_{q,p}(\Omega_{T},wd\mu)}.$
Here in estimates in Theorems 2.2 and 2.3 involving parameters $R_{0},K_{1},K_{2},K_{3}$, we write the explicit dependence on them in order for the constants $\lambda_{0},\varepsilon,M$, and $N$ to be independent of these parameters. Moreover, in the above form, all estimates in Theorems 2.2 and 2.3 are invariant with respect to the parabolic scaling $(t,x)\to(s^{2}t,sx)$, where $s>0$.
@@
    The classes $X_{p}(\mu)$ and $Y_{p}(\mu)$ are special cases of general classes of weights introduced by Muckenhoupt in [47], in which the author obtained a complete characterization of weights admitting certain weighted Hardy’s inequalities on a half-line. The class of $X_{p}(\mu)$ weights is characterized by the validity of the weighted Hardy’s inequality $\|u/x_{d}\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N\|D_{d}u\|_{L_{q,p}(\Omega_{T},wd\mu)}$. See Lemma 6.1 for a precise statement. We use this inequality in the proofs of Theorem 2.2 (ii) and Theorem 2.3 (ii).
@@
    In Theorem 2.3, from the $Y_{p}(\mu)$ condition it also follows that any weak solution $u$ to (1.1) with $u,Du,F,f\in L_{q,p}(\Omega_{T},wd\mu)$ automatically satisfies the conormal boundary condition (1.2). See the end of Section 6 for the proof. This fact is of independent interest, although we do not use it later in the proofs.
@@
Here $A^{ij},C_{0},B^{i},\hat{B}^{i},C:\mathbb{R}^{d}_{+}\to\mathbb{R}^{n\times n}$ and $u,F^{i},F_{0},f:\mathbb{R}^{d}_{+}\to\mathbb{R}^{n}$ are functions of $x\in\mathbb{R}^{d}_{+}$. We say that $u$ is a weak solution to (2.9)-(2.10) if
@@
holds for all $\varphi\in C_{0}^{\infty}(\mathbb{R}^{d};\mathbb{R}^{n})$. We only state the following elliptic analogue of Theorem 2.2. The elliptic analogue of Theorem 2.3 can be stated similarly.
@@
Let $\alpha\in(-1,\infty),p\in(1,\infty),\kappa\in(0,1],K\in[1,\infty),R_{0}\in(0,\infty]$, and $K_{1}\in[0,\infty)$. Let $P=(d,\alpha,p,\kappa,K)$ denote the collection of listed parameters. There exist $\gamma_{0}=\gamma_{0}(P)>0$ and $\lambda_{0}=\lambda_{0}(P)>0$ such that the following assertions hold. Let $w=w(x)\in A_{p}(\mathbb{R}^{d}_{+})$ with $w]_{A_{p}(\mathbb{R}^{d}_{+},\mu)}\leq K$. Suppose that (2.1), (2.2), and Assumption 2.1 $(\gamma_{0},R_{0})$ are satisfied. Also assume that $F_{0}=0$.
@@
    Suppose that $|B|\leq K_{1}x_{d},|\hat{B}|\leq K_{1}x_{d}$, and $|C|\leq K_{1}^{2}x_{d}^{2}$. There exist $M=M(P)>0$ and $N=N(P)>0$ such that for any $F\in L_{p}(\mathbb{R}^{d}_{+},wd\mu;\mathbb{R}^{n})^{d}$, $f\in L_{p}(\mathbb{R}^{d}_{+},wd\mu;\mathbb{R}^{n})$, $u\in W^{1}_{p}(\mathbb{R}^{d}_{+},wd\mu;\mathbb{R}^{n})$, and $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$ satisfying (2.9)-(2.10), it holds that
    $\|Du\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}\leq N\|F\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+N\|f\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}.$
    Moreover, for any $F\in L_{p}(\mathbb{R}^{d}_{+},wd\mu;\mathbb{R}^{n})^{d}$, $f\in L_{p}(\mathbb{R}^{d}_{+},wd\mu;\mathbb{R}^{n})$, and $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$ with $\lambda>0$, there exists a unique solution $u\in W^{1}_{p}(\mathbb{R}^{d}_{+},wd\mu;\mathbb{R}^{n})$ to (2.9)-(2.10).
@@
    Assume, additionally, that $\alpha\in(0,\infty)$, and $w(x)=w_{2}(x^{\prime})w_{3}(x_{d})$, where $w_{2}\in A_{p}(\mathbb{R}^{d-1}),w_{3}\in A_{p}(\mathbb{R}_{+},\mu)\cap X_{p}(\mu)$ with $w_{2}]_{A_{p}(\mathbb{R}^{d-1})}\cdot[w_{3}]_{A_{p}(\mathbb{R}_{+},\mu)}\leq K,[w_{3}]_{X_{p}(\mu)}\leq K_{2},K_{2}\in(0,\infty)$. There exist $\varepsilon=\varepsilon(P)>0,M=M(P)>0$, and $N=N(P)>0$ such that the result of (i) holds under the assumption
    $|B|\leq\max(\varepsilon K_{2}^{-1},K_{1}x_{d}),\quad|\hat{B}|\leq K_{1}x_{d},\quad\text{and}\quad|C|\leq\max(K_{1}K_{2}^{-1}x_{d},K_{1}^{2}x_{d}^{2}).$
    Moreover, in place of ((i)) holds the estimate
    $\|Du\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+K_{2}^{-1}\|u/x_{d}\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}$ $+\sqrt{\lambda}\|u\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}$
    $\leq N\|F\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}+N\|f\|_{L_{p}(\mathbb{R}^{d}_{+},wd\mu)}.$
Theorem 2.5 can be derived from Theorem 2.2 by viewing solutions to elliptic systems as steady state solutions of the corresponding parabolic systems. More precisely, elliptic a-priori estimates above can be derived from estimates (2.5) and (2.6) in Theorem 2.2. See, for example, the proofs of [40, Theorem 2.6] or [17, Theorem 1.2] for details. The existence of solutions can be obtained by following the proof of existence in Theorem 2.2. We therefore omit the proof of Theorem 2.5.